# FusionDTI: Fine-grained Binding Discovery with Token-level Fusion for Drug-Target Interaction

### v3

@@ Abstract
Predicting drug-target interaction (DTI) is critical in the drug discovery process. Despite remarkable advances in recent DTI models through the integration of representations from diverse drug and target encoders, such models often struggle to capture the fine-grained interactions between drugs and protein, i.e. the binding of specific drug atoms (or substructures) and key amino acids of proteins, which is crucial for understanding the binding mechanisms and optimising drug design. To address this issue, this paper introduces a novel model, called FusionDTI, which uses a token-level Fusion module to effectively learn fine-grained information for Drug-Target Interaction. In particular, our FusionDTI model uses the SELFIES representation of drugs to mitigate sequence fragment invalidation and incorporates the structure-aware (SA) vocabulary of target proteins to address the limitation of amino acid sequences in structural information, additionally leveraging pre-trained language models extensively trained on large-scale biomedical datasets as encoders to capture the complex information of drugs and targets. Experiments on three well-known benchmark datasets show that our proposed FusionDTI model achieves the best performance in DTI prediction compared with eight existing state-of-the-art baselines. Furthermore, our case study indicates that FusionDTI could highlight the potential binding sites, enhancing the explainability of the DTI prediction. 111The complete code, datasets and demo are available at: https://github.com/ZhaohanM/FusionDTI..
@@ 1 Introduction
The task of predicting drug-target interactions (DTI) plays a pivotal role in the drug discovery progress, as it helps identify potential therapeutic effects of drugs on biological targets facilitating the development of effective treatments [2]. DTI fundamentally relies on the binding of specific drug atoms (or substructures) and key amino acids of proteins [33]. In particular, each binding site is an interaction between a single amino acid and a single drug atom, which we refer to as a fine-grained interaction. For instance, Figure 1 B demonstrates the interaction between HIV-1 protease and the drug lopinavir. A critical component of this interaction is the formation of a hydrogen bond between a ketone group in lopinavir (represented in the SELFIES [20] notation as [C][=O]) and the side chain of an aspartate residue Asp25 (i.e. Dd) within the protease [5, 7]. Therefore, capturing such fine-grained interaction information during the fusion of drug and target representations is crucial for building effective DTI prediction models [45, 43, 28, 48].
@@
To obtain representations of drugs and targets for the DTI task, some previous studies [21, 26] have used graph neural networks (GNNs) or convolutional neural networks (CNNs) using a fixed-size window, potentially leading to a loss of contextual information, especially when drugs and targets are in a long-term sequence. These models directly concatenate the representations together to make predictions without considering fine-grained interactions. More recently, some computational models [17, 3] employed the fusion module (e.g. Deep Interactive Inference Network (DIIN) [15] and Bilinear Attention Network (BAN) [19]) to obtain fine-grained interaction information and the 3-mer approach that binds three amino acids together as a target binding site to address the lack of structural information in the amino acid sequence. While useful for highlighting possible regions of interaction, these models do not offer the sufficient granularity needed to gauge the specifics of binding sites, as each binding site only contains one residue [33]. Therefore, obtaining contextual representations of drugs and targets and capturing fine-grained interaction information for DTI remains challenging.
@@
To address these challenges, we propose a novel model (called FusionDTI) with a Token-level Fusion (TF) module for an effective learning of fine-grained interactions between drugs and targets. In particular, our FusionDTI model utilises two pre-trained language models (PLMs), namely Saport [34] as the protein encoder that is able to integrate both residue tokens with structure token; and SELFormer [47] as the drug encoder to ensure that each drug is valid and contains structural information. To effectively learn fine-grained information from these contextual representations of drugs and targets, we explore two strategies for the TF module, i.e. Bilinear Attention Network (BAN) [19] and Cross Attention Network (CAN) [22, 38], to find the best approach for integrating the rich contextual embeddings derived from Saport and SELFormer. We conduct a comprehensive performance comparison against eight existing state-of-the-art DTI prediction models. The results show that our proposed model achieves about 6% accuracy improvement over the best baseline on the BindingDB dataset. The main contributions of our study are as follows:
@@
We propose FusionDTI, a novel model that leverages PLMs to encode drug SELFIES and protein residue and structure for rich semantic representations and uses the token-level fusion to obtain fine-grained interaction information between drugs and targets effectively.
@@ 2 Related Work
2.1 Drug-target Interaction Prediction
DTI prediction serves as an important step in the process of drug discovery [10]. Traditional biomedical measurements from wet experiments are reliable but have a notably high cost and time-consuming development cycle, preventing their application on large-scale data [51]. In contrast, identifying high-confidence DTI pairs by computational models markedly narrow down the search scope of drug candidate libraries, and aims to identify drugs most likely to bind to a target. Support vector machine (SVM) [9] and random forest (RF) [16] are two traditional computational models for DTI by concatenating fingerprint ECFP4 [30] and PSC features [6]. Later works focused on representation learning approaches, such as CNNs and GNNs [21, 26]. For example, DeepConv-DTI [21] employed CNNs and a global max-pooling layer to extract local protein sequence patterns. Similarly, GraphDTA [26] used GNNs for drug graph encoding and CNNs for protein sequence encoding. Furthermore, MolTrans [17] introduced an adaptation of the transformer for encoding, further enhanced by a DIIN module [15] to learn fine-grained interactions. Additionally, DrugBAN [3] incorporated a deep BAN [19] framework with domain adaptation to facilitate explicit pairwise fine-grained interaction learning between drugs and targets. Moreover, BioT5 [27] has been proposed as a comprehensive pre-training framework that integrates cross-modelling in biology in the DTI task. More recently, SiamDTI [49] utilises a double-channel network structure to acquire local and global protein information for cross-field DTI prediction. Despite these advances, these models have not proposed an effective way to capture fine-grained interaction information in the DTI.
@@ 2.2 Drug and Protein Representation
For drug molecules, most existing methods represent the input by the Simplified Molecular Input Line Entry System (SMILES) [40, 41]. However, SMILES suffers from numerous problems in terms of validity and robustness, and some valuable information about the drug structure may be lost which may prevent the model from efficiently mining the knowledge hidden in the data reducing the predictive performance of the model [20]. In particular, SMILES fragments are often invalid and inconsistent with the substructural information of the drug. To address the limitations of SMILES, we apply SELFIES [20], a string-based representation that circumvents the issue of robustness and that always generates valid molecular graphs for each character [20].
Regarding proteins, the conventional approach uses amino acid sequences as model inputs [17, 3], overlooking the crucial structural information of the protein. Inspired by the SA vocabulary of SaProt [34], the SaProt enhances inputs by amalgamating each residue from the amino acid sequence with a 3D geometric feature that is obtained by encoding the structure information of the protein using Foldseek [36]. This innovative combination offers richer protein representations through the SA vocabulary, contributing to the discovery of fine-grained interactions. Our proposed model employs SELFIES for drug encoding and uses SaProt encoding for proteins to generate the semantic representations for both drugs and targets.
@@ 2.3 Molecular and Protein Language Models
Molecular language models that train on the large-scale molecular corpus to capture the subtleties of chemical structures and their biological activities have set new standards in encoding chemical compounds achieving meaningful representations [46, 31]. For example, ChemBERTa-2 [1] used RoBERTa-based architectures to capture intricate molecular patterns, significantly enhancing the precision of property prediction. Subsequently, MoLFormer [32] focused on leveraging the self-attention mechanism to interpret the complex, non-linear interactions within molecules, while SELFormer [47] employed SELFIES, ensuring valid and interpretable chemical structures.
Protein language models have revolutionized the way we understand and represent protein sequences, offering richer semantic representations [12, 23, 34]. These models leverage the vast corpus of biological sequence data, learning intricate patterns and features that define the protein functionality and interactions. ProtBERT [12] and ESM [23] applied a transformer architecture to protein sequences, capturing the complex relationships between amino acids. Saport [34] further enhanced this approach by integrating SA vocabularies to provide protein structure information. Furthermore, SaprotHub [35] offers a platform that enables biologists to train, deploy, and share protein models efficiently. Importantly, our FusionDTI is flexible enough to use each of them as a protein encoder.
@@ 3.1 Model Architecture
Given a sequence-based input drug-target pair, the DTI prediction task aims to predict an interaction probability score between the given drug-target pair, which is typically achieved through learning a joint representation space from the given sequence-based inputs. To address the DTI task and effectively capture fine-grained interaction, we proposed a novel model, called FusionDTI, which is a bi-encoder model [24] with a fusion module that fuses the representations of drugs and targets. The overall framework of FusionDTI is illustrated in Figure 1 A. In general, FusionDTI takes sequence-based inputs of drugs and targets, which are encoded into token-level representation vectors by two frozen encoders. Then, a fusion module fuses the representations to capture fine-grained binding information for a final prediction through a prediction head.
Input: The initial inputs of drugs and targets are string-based representations. For protein , the SA vocabulary [34, 36] is employed, where each residue is replaced by one of 441 SA vocabularies that bind an amino acid to a 3D geometric feature to address the lack of structural information in amino acid sequences. For drug , as mentioned in the previous section, we use the SELFIES, which is a formal syntax that always generates valid molecular graphs [20]. We provide the steps and code for obtaining SA and SELFIES sequences in Appendix 6.3.
Encoder: The proposed model contains two frozen encoders: Saport [34] and SELFormer [47], which generate a drug representation and a protein representation separately. It is of note that FusionDTI is flexible enough to easily replace encoders with other advanced PLMs. Furthermore, and are stored in memory for later-stage online training.
Fusion module: In developing FusionDTI, we have investigated two options for the fusion module: BAN and CAN to fuse representations, as indicated in Figure 2. The CAN is utilised to fuse each pair as and , and then concatenate them into one for fine-grained binding information. For BAN, we need to obtain the bilinear attention map and then generate through the bilinear pooling layer.
@@
Since the encoders and the fusion module constitute the key components of our FusionDTI model, we will describe them in detail in the following subsections.
@@ 3.2 Drug and Protein Encoders
Employing sequences with detailed biological functions and structures is a critical step in exploring the fine-grained binding of drugs and targets. For drugs, SMILES is the most commonly used input sequence but suffers from invalid sequence segments and potential loss of structural information [20]. To address the limitations, we transform SMILES into SELFIES, a formal grammar that generates a valid molecular graph for each element [20]. Besides, to address the lack of structural information in the amino acid sequences, we utilise the SA sequence of targets to combine each amino acid with an SA vocabulary by Foldseek [36].
PLMs have shown promising achievements in the biomedical domain leveraging transformers since they pay attention to contextual information and are pre-trained on large-scale biomedical databases. Therefore, we utilise Saport [34] as a protein encoder to encode protein input of both the SA sequence and amino acid sequence. Meanwhile, SELFormer [47] is used as our drug encoder to encode the drug SELFIES input . Then these encoded protein representation and drug representation are further used as inputs for the later fusion module (Subsection 3.3). These rich contextual representations ensure that we can explore the fine-grained binding information effectively. To further justify this, we also compare our encoders with other existing protein language models (such as ESM-2 [23]) and molecular language models (such as MoLFormer [32] and ChemBERTa-2 [1]), and the results can be found in Section 4.7.
@@ 3.3 Fusion Module
In order to capture the fine-grained binding information between a drug and a target, our FusionDTI model applies a fusion module to learn token-level interactions between the token representations of drugs and targets encoded by their respective encoders. As shown in Figure 2, two fusion modules inspired by the recent literature [3, 44] are investigated to fuse representations: the Bilinear Attention Network [19] and the Cross Attention Network [22, 38].
@@ 3.3.1 Bilinear Attention Network (BAN)
Motivated by DrugBAN [3], our model considers BAN [19] as an option of the fusion module to learn pairwise fine-grained interactions between drug and target , denoted as FusionDTI-BAN. For BAN as indicated in Figure 2, bilinear attention maps are obtained by a bilinear interaction modelling to capture pairwise weights in step 1, and then the bilinear pooling layer to extract a joint representation . The equation for BAN is shown below:
@@
Here, is a fixed all-ones vector, is a learnable weight vector and denotes the Hadamard product. In this way, pairwise interactions contribute sub-structural pairs to the prediction.
@@ 3.3.2 Cross Attention Network (CAN)
Inspired by ProST [44], we also consider CAN as our fusion module to learn fine-grained interaction information of drugs and targets. We denote our FusionDTI model that uses a CAN fusion module as FusionDTI-CAN. By processing and separately, the fused drug and target representations are obtained. To synthesise the fine-grained joint representation , we employ a pooling aggregation strategy for both and independently and then concatenate them as shown in Figure 2. The process is delineated by the following equation:
@@
In summary, our CAN module combines multi-head, self-attention and cross-attention mechanisms to capture dependencies within individual sequences and between different sequences for a more nuanced understanding of interactions. In the results of Sections 4.3 and 4.6, we analyse and compare these two fusion strategies and different fusion scales in detail.
@@ 4 Experimental Setup and Results
4.1 Datasets and Baselines
Three public DTI datasets, namely BindingDB [14], BioSNAP [50] and Human [25, 8], are used for evaluation, where each dataset is split into training, validation, and test sets with a 7:1:2 ratio using two different splitting strategies: in-domain and cross-domain. For the in-domain split, the datasets are randomly divided. For the cross-domain setting, the datasets are split such that the drugs and targets in the test set do not overlap with those in the training set, making it a more challenging scenario where models must generalise to novel drug-target interactions. Since DTI is a binary classification task, we use AUROC (area under the receiver operating characteristic curve) [3, 17] and AUPRC (area under the precision-call curve) [21, 26] as the major metrics to evaluate a model’s performance.
We compare FusionDTI with eight baseline models in the DTI prediction task. These models include two traditional machine learning methods such as SVM [9] and Random Forest (RF) [16], as well as five deep learning methods including DeepConv-DTI [21], GraphDTA [26], MolTrans [17], DrugBAN [3] and SiamDTI [49]. The latter five models employ the same two-stage process whereby the drug and target features are initially extracted by specialised encoders before being integrated for prediction. In addition, we also include the BioT5 [27] model, which is a biomedical pre-trained language model that could directly predict the DTI. Further details regarding the datasets, baseline models, and the methodology for generating drug SELFIES and protein SA sequences are provided in Appendix 6.3.
4.2 Effectiveness Evaluation for DTI Prediction
@@
| FusionDTI-BAN | .659.002 | .663.002 | .633.003 | .784.002 | .790.003 | .723.002 | .721.002 | .756.001 |
| FusionDTI-CAN | .675.005 | .676.012 | .649.005 | .801.037 | .803.032 | .748.021 | .766.017 | .734.012 |

### v4

@@ Abstract
Predicting drug-target interaction (DTI) is critical in the drug discovery process. Despite remarkable advances in recent DTI models through the integration of representations from diverse drug and target encoders, such models often struggle to capture the fine-grained interactions between drugs and protein, i.e. the binding of specific drug atoms (or substructures) and key amino acids of proteins, which is crucial for understanding the binding mechanisms and optimising drug design. To address this issue, this paper introduces a novel model, called FusionDTI, which uses a token-level Fusion module to effectively learn fine-grained information for Drug-Target Interaction. In particular, our FusionDTI model uses the SELFIES representation of drugs to mitigate sequence fragment invalidation and incorporates the structure-aware (SA) vocabulary of target proteins to address the limitation of amino acid sequences in structural information, additionally leveraging pre-trained language models extensively trained on large-scale biomedical datasets as encoders to capture the complex information of drugs and targets. Experiments on three well-known benchmark datasets show that our proposed FusionDTI model achieves the best performance in DTI prediction compared with eight existing state-of-the-art baselines. Furthermore, our case study indicates that FusionDTI could highlight the potential binding sites, enhancing the explainability of the DTI prediction.
FusionDTI: Fine-grained Binding Discovery with Token-level Fusion for Drug-Target Interaction
Zhaohan Meng1,
Zaiqiao Meng1††thanks: Corresponding author,
Ke Yuan2,3 &
Iadh Ounis1
1School of Computing Science 2School of Cancer Sciences
3Cancer Research UK Scotland Institute
University of Glasgow, United Kingdom
{z.meng.3}@research.gla.ac.uk
{zaiqiao.meng, ke.yuan, iadh.ounis}@glasgow.ac.uk
Source Code
Demo Space
@@ 1 Introduction
The task of predicting drug-target interactions (DTI) plays a pivotal role in the drug discovery progress, as it helps identify potential therapeutic effects of drugs on biological targets facilitating the development of effective treatments (Askr et al., 2023). DTI fundamentally relies on the binding of specific drug atoms (or substructures) and key amino acids of proteins (Schenone et al., 2013). In particular, each binding site is an interaction between a single amino acid and a single drug atom, which we refer to as a fine-grained interaction. For instance, Figure 1 B demonstrates the interaction between HIV-1 protease and the drug lopinavir. A critical component of this interaction is the formation of a hydrogen bond between a ketone group in lopinavir (represented in the SELFIES (Krenn et al., 2022) notation as [C][=O]) and the side chain of an aspartate residue Asp25 (i.e. Dd) within the protease (Brik and Wong, 2003; Chandwani and Shuter, 2008). Therefore, capturing such fine-grained interaction information during the fusion of drug and target representations is crucial for building effective DTI prediction models (Wu et al., 2022; Peng et al., 2024; Zeng et al., 2024).
@@
To obtain representations of drugs and targets for the DTI task, some previous studies (Lee et al., 2019; Nguyen et al., 2021) have used graph neural networks (GNNs) or convolutional neural networks (CNNs) using a fixed-size window, potentially leading to a loss of contextual information, especially when drugs and targets are in a long-term sequence. These models directly concatenate the representations together to make predictions without considering fine-grained interactions. More recently, some computational models (Huang et al., 2021; Bai et al., 2023) employed the fusion module (e.g. Deep Interactive Inference Network (DIIN) (Gong et al., 2018) and Bilinear Attention Network (BAN) (Kim et al., 2018)) to obtain fine-grained interaction information and the 3-mer approach that binds three amino acids together as a target binding site to address the lack of structural information in the amino acid sequence. While useful for highlighting possible regions of interaction, these models do not offer the sufficient granularity needed to gauge the specifics of binding sites, as each binding site only contains one residue (Schenone et al., 2013). Therefore, obtaining contextual representations of drugs and targets and capturing fine-grained interaction information for DTI remains challenging.
@@
To address these challenges, we propose a novel model (called FusionDTI) with a Token-level Fusion (TF) module for an effective learning of fine-grained interactions between drugs and targets. In particular, our FusionDTI model utilises two pre-trained language models (PLMs), namely Saport (Su et al., 2023) as the protein encoder that is able to integrate both residue tokens with structure token; and SELFormer (Yüksel et al., 2023) as the drug encoder to ensure that each drug is valid and contains structural information. To effectively learn fine-grained information from these contextual representations of drugs and targets, we explore two strategies for the TF module, i.e. Bilinear Attention Network (BAN) (Kim et al., 2018) and Cross Attention Network (CAN) (Li et al., 2021; Vaswani et al., 2017), to find the best approach for integrating the rich contextual embeddings derived from Saport and SELFormer. We conduct a comprehensive performance comparison against eight existing state-of-the-art DTI prediction models. The results show that our proposed model achieves about 6% accuracy improvement over the best baseline on the BindingDB dataset. The main contributions of our study are as follows:
@@
We propose FusionDTI, a novel model that leverages PLMs to encode drug SELFIES, as well as protein residues and structures for rich semantic representations and uses the token-level fusion to capture fine-grained interaction between drugs and targets effectively.
@@ 2.1 Drug and Protein Representation
For drug molecules, most existing methods represent the input by the Simplified Molecular Input Line Entry System (SMILES) (Weininger, 1988; Weininger et al., 1989). However, SMILES suffers from numerous problems in terms of validity and robustness, and some valuable information about the drug structure may be lost which may prevent the model from efficiently mining the knowledge hidden in the data (Krenn et al., 2022). To address the limitations of SMILES, we apply SELFIES, a string-based representation that circumvents the issue of robustness and that always generates valid molecular graphs for each character.
Regarding proteins, the conventional approach uses amino acid sequences as model inputs (Huang et al., 2021; Bai et al., 2023), overlooking the crucial structural information of the protein. Inspired by the SA vocabulary of SaProt (Su et al., 2023), the SaProt enhances inputs by amalgamating each residue of the amino acid sequence with a 3D geometric feature that is obtained by encoding protein structure information using Foldseek (Van Kempen et al., 2024). This innovative combination offers richer protein representations through the SA vocabulary, contributing to the discovery of fine-grained interactions.
@@ 2.2 Molecular and Protein Language Models
Molecular language models trained on the large-scale molecular corpus capture the subtleties of chemical structures and their biological activities, setting new standards in the encoding of chemical compounds achieving meaningful representations Ying et al. (2021); Rong et al. (2020). For example, MoLFormer (Ross et al., 2022) focused on leveraging the self-attention mechanism to interpret the complex, non-linear interactions within molecules, while SELFormer (Yüksel et al., 2023) employed SELFIES, ensuring valid and interpretable chemical structures.
Protein language models have revolutionized the way we understand and represent protein sequences, learning intricate patterns and features that define the protein functionality and interactions. ProtBERT (Elnaggar et al., 2021) and ESM (Lin et al., 2023) applied a transformer architecture to protein sequences, capturing the complex relationships between amino acids. Saport (Su et al., 2023, 2024) further enhanced this approach by integrating SA vocabularies to provide protein structure information.
@@ 3.1 Model Architecture
Given a sequence-based input drug-target pair, the DTI prediction task aims to predict an interaction probability score between the given drug-target pair, which is typically achieved through learning a joint representation space from the given sequence-based inputs. To address the DTI task and effectively capture fine-grained interaction, we proposed a novel model, called FusionDTI, which is a bi-encoder model Liu et al. (2021) with a fusion module that fuses the representations of drugs and targets. The overall framework of FusionDTI is illustrated in Figure 1 A. In general, FusionDTI takes sequence-based inputs of drugs and targets, which are encoded into token-level representation vectors by two frozen encoders. Then, a fusion module fuses the representations to capture fine-grained binding information for a final prediction through a prediction head.
Input: The initial inputs of drugs and targets are string-based representations. For protein , the SA vocabulary (Su et al., 2023; Van Kempen et al., 2024) is employed, where each residue is replaced by one of 441 SA vocabularies that bind an amino acid to a 3D geometric feature to address the lack of structural information in amino acid sequences. For drug , as mentioned in the previous section, we use the SELFIES, which is a formal syntax that always generates valid molecular graphs (Krenn et al., 2022). We provide the steps and code to obtain SA and SELFIES in Appendix A.3.
Encoder: The proposed model contains two frozen encoders: Saport (Su et al., 2023) and SELFormer (Yüksel et al., 2023), which generate a drug representation and a protein representation separately. It is of note that FusionDTI is flexible enough to easily replace encoders with other PLMs or address SELFIES or SA representations that are unavailable. Furthermore, and are stored in memory for later-stage online training.
Fusion module: In developing FusionDTI, we have investigated two options for the fusion module: BAN and CAN to fuse representations, as indicated in Figure 2. The CAN is utilised to fuse each pair as and , and then concatenate them into one for fine-grained binding information. For BAN, we need to obtain bilinear attention maps and generate through the bilinear pooling layer.
@@
Since the encoders and the fusion module constitute the key components of our FusionDTI model, we will describe them in detail in the following.
@@ 3.2 Drug and Protein Encoders
Employing sequences with detailed biological functions and structures is a critical step in exploring the fine-grained binding of drugs and targets. For drugs, SMILES is the most commonly used input sequence but suffers from invalid sequence segments and potential loss of structural information (Krenn et al., 2022). To address the limitations, we transform SMILES into SELFIES, a formal grammar that generates a valid molecular graph for each element (Krenn et al., 2022). Besides, to address the lack of structural information in the amino acid sequences, we utilise the SA sequence of targets to combine each amino acid with an SA vocabulary by Foldseek (Van Kempen et al., 2024).
PLMs have shown promising achievements in the biomedical domain leveraging transformers since they pay attention to contextual information and are pre-trained on large-scale biomedical databases. Therefore, we utilise Saport (Su et al., 2023) as a protein encoder to encode protein input of both the SA sequence and amino acid sequence. Meanwhile, SELFormer (Yüksel et al., 2023) is used as our drug encoder to encode the drug SELFIES input . Then these encoded protein representation and drug representation are further used as inputs for the later fusion module (Subsection 3.3). These rich contextual representations ensure that we can explore the fine-grained binding information effectively. To further justify this, we also compare our encoders with other existing protein language models (such as ESM-2 (Lin et al., 2023)) and molecular language models (such as MoLFormer (Ross et al., 2022) and ChemBERTa-2 (Ahmad et al., 2022)), and the results can be found in Appendix A.6.
@@ 3.3 Fusion Module
In order to capture the fine-grained binding information between a drug and a target, our FusionDTI model applies a fusion module to learn token-level interactions between the token representations of drugs and targets encoded by their respective encoders. As shown in Figure 2, two fusion modules are investigated to fuse representations: the Bilinear Attention Network (Kim et al., 2018) and the Cross Attention Network (Vaswani et al., 2017).
@@ 3.3.1 Bilinear Attention Network (BAN)
Motivated by DrugBAN (Bai et al., 2023), our model considers BAN (Kim et al., 2018) as an option to learn pairwise fine-grained interactions between drug and target , denoted as FusionDTI-BAN. For BAN as indicated in Figure 2, bilinear attention maps are obtained by a bilinear interaction modelling to capture pairwise weights in step 1, and then the bilinear pooling layer to extract a joint representation . The equation of BAN is shown below:
@@
Here, is a fixed all-ones vector, is a learnable weight vector and denotes the Hadamard product. In this way, pairwise interactions contribute sub-structural pairs to predictions.
@@ 3.3.2 Cross Attention Network (CAN)
Inspired by ProST (Xu et al., 2023), we also consider CAN as our fusion module to learn fine-grained interaction information of drugs and targets. We denote our FusionDTI model that uses a CAN fusion module as FusionDTI-CAN. By processing and separately, the fused drug and target representations are obtained. To synthesise the fine-grained joint representation , we employ a pooling aggregation strategy for both and independently and then concatenate them as shown in Figure 2. The process is described by the following equation:
@@
In summary, our CAN module combines multi-head, self-attention and cross-attention mechanisms to capture dependencies within individual sequences and between different sequences for a more nuanced understanding of interactions. In the results of Sections 4.3 and 4.5, we analyse and compare these two fusion strategies and different fusion scales in detail.
@@ 4 Experimental Setup and Results
| FusionDTI-BAN | .659.002 | .663.002 | .633.003 | .784.002 | .790.003 | .723.002 | .721.002 | .734.001 |
| FusionDTI-CAN | .681.005 | .680.012 | .652.005 | .801.037 | .803.032 | .748.021 | .766.017 | .756.012 |
4.1 Datasets and Baselines
Three public DTI datasets, namely BindingDB (Gilson et al., 2016), BioSNAP (Zitnik et al., 2018) and Human (Liu et al., 2015; Chen et al., 2020), are used for evaluation, where each dataset is split into training, validation, and test sets with a 7:1:2 ratio using two different splitting strategies: in-domain and cross-domain. For the in-domain split, the datasets are randomly divided. For the cross-domain setting, the datasets are split such that the drugs and targets in the test set do not overlap with those in the training set, making it a more challenging scenario where models must generalise to novel drug-target interactions. Since DTI is a binary classification task, we use AUROC (Bai et al., 2023; Huang et al., 2021) and AUPRC (Nguyen et al., 2021) as the major metrics to evaluate models’ performance. In Appendix A.10, we report other evaluation metrics, including F1-score, Sensitivity, Specificity, and Matthews Correlation Coefficient (MCC) to provide a more comprehensive assessment.
We compare FusionDTI with eight baseline models in the DTI prediction task. These models include two traditional machine learning methods such as SVM (Cortes and Vapnik, 1995) and Random Forest (RF) (Ho, 1995), as well as five deep learning methods including DeepConv-DTI (Lee et al., 2019), GraphDTA (Nguyen et al., 2021), MolTrans (Huang et al., 2021), DrugBAN (Bai et al., 2023) and SiamDTI (Zhang et al., 2024). In addition, we also include the BioT5 (Pei et al., 2023) model, which is a biomedical pre-trained language model that could directly predict the DTI.
Furthermore, results on three additional benchmark datasets (DAVIS (Davis et al., 2011), KIBA (Tang et al., 2014), and DUD-E (Mysinger et al., 2012)) are reported, with comparisons to 8 task-specific baselines (Nga et al., 2025; Li et al., 2025). Further details regarding the datasets, baseline models, and the methodology for generating drug SELFIES and protein SA sequences are provided in Appendix A.3.
4.2 Evaluation of DTI Prediction